equation of a stretched string is y=ae^-(bx+ct).The direction of propagation of wave is?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
 Y(x,t)=f(x+or-vt)

Let,bx+ct=C
=>x=C/b –ct/b
=>dx/dt=-c/b
=>v=-c/b

V is +ve or –ve will entirely depend on the values of b and c.
In the event that it is negative then the condition will be (x,t)=f(x-vt) and the wave is along these lines spreading in +ve heading .
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)